Solve dy/dx= 4x² (y-x)² + if y /x is a particular solution.


1
Expert's answer
2021-08-18T08:00:01-0400

"\frac{dy}{dx}= 4x² (y-x)²\\
dy= 4x² (y-x)²dx\\
dy- 4x² (y-x)²dx=0\\"

We know that "\frac{\partial F}{\partial x}=\left(4x^2(y-x)^2\right)"

"\implies \int\frac{\partial F }{\partial x}=\int\left(4x^2(y-x)^2\right)\\
\mathrm{Take\:the\:constant\:out}:\quad \int a\cdot f\left(x\right)dx=a\cdot \int f\left(x\right)dx\\
=4\cdot \int \left(y-x\right)^2x^2dx\\
4\cdot \int \:y^2x^2-2yx^3+x^4dx\\
\mathrm{Apply\:the\:Sum\:Rule}:\quad \int f\left(x\right)\pm g\left(x\right)dx=\int f\left(x\right)dx\pm \int g\left(x\right)dx\\
=4\left(\int \:y^2x^2dx-\int \:2yx^3dx+\int \:x^4dx\right)\\
=4\left(\frac{y^2x^3}{3}-\frac{yx^4}{2}+\frac{x^5}{5}\right)+f'(y)\\
\frac{d}{dy}\left(4\left(\frac{y^2x^3}{3}-\frac{yx^4}{2}+\frac{x^5}{5}\right)\right)+ f'(y)\\
=4\frac{d}{dy}\left(\frac{y^2x^3}{3}-\frac{yx^4}{2}+\frac{x^5}{5}\right)+\int f'(y)\\\\
=4\left(\frac{d}{dy}\left(\frac{y^2x^3}{3}\right)-\frac{d}{dy}\left(\frac{yx^4}{2}\right)+\frac{d}{dy}\left(\frac{x^5}{5}\right)\right)+\int f'(y)\\\\
=4\left(-\frac{x^4}{2}+\frac{2yx^3}{3}\right)+\int f(y)\\
F= 4\left(-\frac{x^4}{2}+\frac{2yx^3}{3}\right)+y\\"